# On Edge Coloring of Multigraphs

### v4

@@ 3 Extented Tashkinov Tree (ETT)
Since we did not stop at $\mu_{l-1}$, by the construction algorithm of $\mu_{l}$, we have $\alpha_{i}\notin\overline{\mu}_{l-1}(x_{l})$ and thus $\alpha_{i}\neq\beta_{l}$. Moreover, $\beta_{l-1}\neq\alpha_{i}$ by (6) (when $l=1$) and the construction algorithm of $\mu_{l-1}$ (when $l\geq 2$). So following (6) on $\varphi_{i}=\mu_{0}$, (10), the elementariness of $T^{l-1}_{\mu}$ by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$), $(A_{k})$ for all $11\leq k\leq l-$ 1 (when $l\geq 2$), $(C_{l-1})$, $(F_{l-1})$ (when $l\geq 2$) and $(c_{i})$ (when $l=1$), we have
(13) $\beta_{l},\beta_{l-1}\pi,\alpha_{i},\eta,\gamma_{h}$ (when $h\geq 1$) and $\tau$ are all different, $\pi,\tau\notin S_{h}$ and $\beta_{l}\notin\mu_{l}(P\cup P_{ex}\cup D)$.
@@
Note that $T^{l-1}$ is an ETT with the same $(j-1)$-rungs under $\mu_{l-1}$ as $T_{j}$ under $\varphi$ by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$). Clearly $\mu^{\prime}$ is $(T^{l-1}_{\mu},D_{j-1},\mu_{l-1})$-stable as $T^{l-1}$ is closed for both $\beta_{i}$ and $\tau$ under $\mu_{l-1}$ and $\mu^{\prime}=\mu_{l-1}(G-T^{l-1}_{\mu},\beta_{l},\tau)$. By $(F_{l-1})$ (when $l\geq 2$) and (10) (when $l=1$), we see that $P_{y_{l}}(\beta_{l},\beta_{l-1},\mu^{\prime})\cap T^{l-1}_{\mu}=\{y_{l}\}$. So $\mu_{l}^{\prime}$ is $(T^{l-1}_{\mu}(y_{l})-y_{l},D_{j-1},\mu_{l-1})$-stable. As $T^{l-1}_{\mu}$ is closed for both $\beta_{l}$ and $\tau$ under $\mu_{l-1}$, all the $(\tau,\beta_{l})$-chains intersect $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$ are contained in $G[V(T^{l-1}_{\mu})]$ except for $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ which contains $y_{l}$ who is the only $(\tau,\beta_{l})$-exit of $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$, and all vertices in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$ are contained in $T^{l-1}_{\mu}$. Moreover, since $\beta_{l}\notin\{\tau,\beta_{l-1}\}$ by (13), $y_{l-1}\prec y_{l}$ by (12), and $T^{l}_{\mu}$ is elementary by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$), we have $\beta_{l}\notin(\overline{\mu}_{l-1}(T^{l-1}_{\mu}(y_{l}))\cup\{\beta_{l-1}\}$. So $\beta_{l}\notin\mu_{l-1}(P\cup P_{ex}\cup D\cup P_{1})$ by $(C_{l-1})$ (when $l\geq 2$) and $(c_{i})$ (when $l=1$). Thus $\tau$ is not used on edges of $P\cup P_{ex}\cup D\cup P_{1}-G[V(T^{l-1}_{\mu})]$ and $\beta_{l}$ is not used on edges of $(P\cup P_{ex}\cup D\cup P_{1})\cap G[V(T^{l-1}_{\mu})]$ under $\mu^{\prime}$ and $\mu_{l}^{\prime}$. Therefore, since all vertices in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$ are contained in $T^{l-1}_{\mu}$, none of the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ under $\mu_{l}^{\prime}$ is in the subpath of $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$. Consequently, we have $(E_{l})$ as we switch all the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ not in $P_{r}(\pi,\beta_{l},\mu_{l}^{\prime})$ to $\tau$ by Kempe changes, and none of these $(\tau,\beta_{l})$-chains intersects $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$. Furthermore,
@@
So we may assume that $l=1$. So $\beta_{0}=\eta=\delta_{h}$. In this case, $P_{y_{1}}(\beta_{1},\delta_{h},\mu^{\prime})$ is the same path as the $(\tau,\delta_{h})$-exit path of $T_{i}$, which is the subpath $P_{1}(y_{1}r^{\prime})$ of $P_{1}(wr^{\prime})$. Since $T_{n+1}$ is $RE$ finished under $\varphi$, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{n+1})]$. Since $P_{1}(wr^{\prime})$ does not intersect $T_{n+1}$ by assumption (iv) of $P$, $P_{1}(y_{1}r^{\prime})$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi$. Because all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are the same as under $\varphi_{i}$ by $(a_{k})$ for all $1\leq k\leq i$, we see that $P_{1}(y_{1}r^{\prime})$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi_{i}=\mu_{0}$. Thus all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\mu_{0}$ are the same as under $\mu_{1}^{\prime}$ by the construction of $\mu_{1}^{\prime}$. Since $\beta_{1}\neq\delta_{h}$ ($\beta_{0}=\delta_{h}$) and $\tau\notin S_{h}$ (see (13)), we see that all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\mu_{1}$ are the same as under $\mu_{1}^{\prime}$. Thus we have $(A_{1})$ by (14).
By (A1), (A2), $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$) on $T_{h}$ under $\mu_{l-1}$, we have
@@
By $(A_{l})$ and (15), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\mu_{l}$ as under $\mu_{l-1}$ if $h\geq 1$. Recall that $y_{l}\notin T_{j-1}$ by (12). Since $\mu_{l}$ is $(T^{l-1}_{\mu}(y_{l})-y_{l},D_{j-1},\mu_{l-1})$-stable, $y_{l}\notin T_{j-1}$, and $T^{l-1}_{\mu}$ has the same rung under $\mu_{l-1}$ as $T_{j}$ under $\varphi$ by $(B_{j-1})$ (when $j\geq 2$) and $(b_{i})$ (when $j=1$), we see that $\mu_{l}$ is also $(T_{j-1},D_{j-1},\mu_{l-1})$-stable, and thus is a $T_{j-1}+f_{j-1}$ mod $\mu_{l-1}$ coloring by (A2.5). Therefore, following Corollary 3.1, $T^{l-1}_{\mu}(y_{l})$ is an ETT with the same $(j-1)$-rungs under $\mu_{l}$ as $T^{l-1}_{\mu}$ under $\mu_{l-1}$. Since $T^{i}$ is a closure of $T^{l-1}_{\mu}(y_{l})$ under $\mu_{i}$ and $j-1\leq n$, we have $(B_{l})$ by (A1) on $T^{l}_{\mu}$, $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$).
@@
We claim that $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$. Otherwise, by $(C_{l})$, we have that $V(P_{1})\subseteq V(T^{l}_{\mu})$ by TAA. So $w\in V(T^{l}_{\mu})$. Because $T^{l}_{\mu}$ is elementary by $(B_{l})$, we have $\alpha_{i}\in\overline{\mu}_{l}(w)$ by $(D_{l})$. So $u$ and $v$, the two ends of $P$ are also in $T^{l}_{\mu}$ by $(C_{l})$ and TAA. However, this is a contradiction to $(D_{l})$ and the elementariness of $T^{l}_{\mu}$. So $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$, and therefore $T^{l}_{\mu}$ must have a $(\tau,\beta_{l})$-exit vertex $y_{l+1}$ in $P_{r}(\tau,\beta_{l},\mu_{l})$ by the elementariness of $T^{l}_{\mu}$. This proves the first half of $(F_{l})$.
We then claim that $P_{r}(\tau,\beta_{l},\mu_{l})$ contains a vertex in $T^{l}_{\mu}$ that is after $y_{l}$ along the direction from $r$. Indeed, if $w\notin T^{l}_{\mu}$, then the two edge disjoint paths $P_{1}(rw)$ and $P_{1}(wy_{1})$ provide us two $\beta_{l}$ colored boundary edges of $T^{l}_{\mu}$ which are also contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$ by $(E_{l})$ and $(C_{l})$, and one of them provide a vertex in $T^{l}_{\mu}$ that is after $y_{l}$ along the direction from $r$. So we assume that $w\in T^{l}_{\mu}$. Then $\alpha_{i}\in\overline{\mu}_{l}(w)$ by $(D_{l})$ and the elementariness of $T^{l}_{\mu}$ followed from $(B_{l})$. If one of $u$ and $v$, say $v$, is contained in $T^{l}_{\mu}$, then $\beta_{i}\in\overline{\mu}_{l}(v)$ by the elementariness of $T^{l}_{\mu}$ and $(D_{l})$. So both $u$ and $v$ are contained in $T^{l}_{\mu}$ by TAA and $(C_{l})$, a contradiction to $(D_{l})$ and the elementariness of $T^{l}_{\mu}$. Thus none of $u$ and $v$ is contained in $T^{l}_{\mu}$, and in this case $P_{2}$ and $P_{3}$ (see (3)) provide us two $\beta_{l}$ colored boundary edges of $T^{l}_{\mu}$ which are also contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$ by $(E_{l})$ and $(C_{l})$. Similarly as earlier, we have as claimed.
@@
With the six claims above and the construction algorithm of $\mu_{l}$ earlier, we see that we can always recursively construct $\mu_{l}$ with a new vertex $y_{l+1}$ with $x_{i}=y_{0}\prec y_{1},...,y_{l+1}$ along $T^{l}_{\mu}$ by (12), unless $\alpha_{i}\in\overline{\mu}_{l}(T^{l}_{\mu}(y_{l+1}))$. But since we have a finite graph, at a certain point we have to stop at some $\mu_{l}$ with $\alpha_{i}\in\overline{\mu}_{l}(T^{l}_{\mu}(y_{l+1}))$. For convenience, we still denote the corresponding coloring by $\mu_{l}$ and use index $l$ for everything we established earlier. Note that $l=0$ may happen.
@@
By $(F_{l})$ (when $l\geq 1$) and (10) (when $l=0$), we see that $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$. So following the elementariness of $T^{l}_{\mu}$, $(D_{l})$ (when $l\geq 1$) and $(d_{i})$ (when $l=0$), none of $u$ and $v$ is contained in $T^{l}_{\mu}$. Thus following TAA, (16), $(C_{l})$ (when $l\geq 1$) and $(c_{i})$ (when $l=0$), the paths $P_{2}$ and $P_{3}$ (see (3)) contain two $\beta_{l}$ colored boundary edges $f$ and $g$ of $T^{l}_{\mu}$, where $f\in P_{2}$ is incident with $u^{\prime}\in V(T^{l}_{\mu})$ and $g\in P_{3}$ is incident with $v^{\prime}\in V(T^{l}_{\mu})$, such that $P_{2}(u^{\prime}u)$ and $P_{3}(v^{\prime}v)$ do not share edge with $G[V(T^{l}_{\mu})]$. Note that all the $\beta_{l}$-colored edges of $P_{2}\cup P_{3}$ are contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ by $(E_{l})$ (when $l\geq 1$), $(c_{i})$ and assumption (iv) of $P$ (when $l=0$). Therefore, since $y_{l+1}$ is the $(\tau,\beta_{l})$-exit vertex of $T^{l}_{\mu}$ in $P_{r}(\tau,\beta_{l},\mu_{l})$, we see that both $u^{\prime}$ and $v^{\prime}$ are in the subpath of $P_{r}(\tau,\beta_{l},\mu_{l})$ from $r$ to $y_{l+1}$. We then let $u^{*}$ (and $v^{*}$) be the nearest vertex in $P_{2}$ (and $P_{3}$) to $u$ (and $v$) shared by the subpath of $P_{r}(\tau,\beta_{l},\mu_{l})$ from $r$ to $y_{l+1}$ and $P_{2}(u^{\prime}u)$ (and $P_{3}(v^{\prime}v)$), respectively. Without loss of the generality, we may assume $u^{*}$ is before $v^{*}$ along $P_{r}(\tau,\beta_{l},\mu_{l})=P_{4}$ from $r$. Now $P_{5}=P_{4}(ru^{*})\cup P_{2}(u^{*}u)$ and $P_{6}=P_{4}(y_{l+1}v^{*})\cup P_{3}(v^{*}v)$ are two edge disjoint paths from $r$ to $u$ and from $y_{l+1}$ to $v$, respectively. Moreover, by (13), (16), $(C_{l})$ (when $l\geq 1$) and $(c_{i})$ (when $l=0$), we have
(17) $\mu_{l}(P_{5}\cup P_{6})\subset(\overline{\mu}_{l}(T^{l}_{\mu}(y_{l}))\cup\{\beta_{l}\}-\{\alpha_{i}\})$.
@@
Finally, we consider the coloring $\phi=\mu_{l}/(G-T^{l}_{\mu},\beta_{l+1},\tau)$, where $\beta_{l+1}$ is an arbitrary color in $\overline{\mu}_{l}(y_{l+1})$. Note that $\beta_{l+1}=\alpha_{i}$ may happen. It is now routine to check that $T^{l}_{\mu}$ remains an ETT with the same $j-1$-rungs under $\phi$ as $T_{j}$ under $\varphi$. Moreover, $P_{y_{l}}(\tau,\beta_{l+1},\phi)$ is the $(\tau,\beta_{l+1})$-exit path of $T^{l}_{\mu}$. By (17), no matter whether $\beta_{l+1}=\alpha_{i}$ or not, by the elementariness of $T^{l}_{\mu}$ under $\mu_{l}$ and $\phi$, we see that $\beta_{l+1}\notin\mu_{l}(P_{5}\cup P_{6})$. Thus $\phi(P_{5}\cup P_{6})\subset(\overline{\phi}(T^{l}_{\mu}(y_{l+1}))\cup\{\beta_{l}\})$, $\beta_{l+1}\notin\phi(P_{5}\cup P_{6})\cap G[V(T^{l}_{\mu})])$, and $\tau\notin\phi(P_{5}\cup P_{6})-G[V(T^{l}_{\mu})])$. In addition, it is also routine to check that $\tau\notin(\{\beta_{l+1}\}\cup S_{h})$. In the case that $\beta_{l+1}=\delta_{h}$, we see that all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T^{l}_{\mu})]$ under $\phi$, so $P_{y_{l}}(\tau,\beta_{l+1},\phi)$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\phi$. In addition, it is also routine to check that $\tau\notin(\{\beta_{l+1}\}\cup S_{h})$ and $(ii)$ in Lemma 4.5 holds for $P_{5}$ and $P_{6}$. So by applying Lemma 4.5 with $P_{y_{l}}(\tau,\beta_{l+1},\phi)$ in place of $P_{ex}$, $r$ in place of $z$ and $\tau$ in place of $\pi$, we have a contradiction between $(II)$ and $(III)$, because $P_{5}$ and $P_{6}$ have different ends $u$ and $v$.

### v5

@@ 3 Extented Tashkinov Tree (ETT)
Since we did not stop at $\mu_{l-1}$, by the construction algorithm of $\mu_{l}$, we have $\alpha_{i}\notin\overline{\mu}_{l-1}(x_{l})$ and thus $\alpha_{i}\neq\beta_{l}$. Moreover, $\beta_{l-1}\neq\alpha_{i}$ by (6) (when $l=1$) and the construction algorithm of $\mu_{l-1}$ (when $l\geq 2$). So following (6) on $\varphi_{i}=\mu_{0}$, (10), the elementariness of $T^{l-1}_{\mu}$ by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$), $(A_{k})$ for all $1\leq k\leq l-1$ (when $l\geq 2$), $(C_{l-1})$, $(F_{l-1})$ (when $l\geq 2$) and $(c_{i})$ (when $l=1$), we have
(13) $\beta_{l},\beta_{l-1}\pi,\alpha_{i},\eta,\gamma_{h}$ (when $h\geq 1$) and $\tau$ are all different, and $\pi,\tau\notin S_{h}$.
@@
Note that $T^{l-1}$ is an ETT with the same (j1)(j-1)-rungs under $\mu_{l-1}$ as $T_{j}$ under $\varphi$ by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$). Clearly $\mu^{\prime}$ is $(T^{l-1}_{\mu},D_{j-1},\mu_{l-1})$-stable as $T^{l-1}$ is closed for both $\beta_{i}$ and $\tau$ under $\mu_{l-1}$ and $\mu^{\prime}=\mu_{l-1}/(G-T^{l-1}_{\mu},\beta_{l},\tau)$. By $(F_{l-1})$ (when $l\geq 2$) and (10) (when $l=1$), we see that $P_{y_{l}}(\beta_{l},\beta_{l-1},\mu^{\prime})\cap T^{l-1}_{\mu}=\{y_{l}\}$. So $\mu_{l}^{\prime}$ is $(T^{l-1}_{\mu}(y_{l})-y_{l},D_{j-1},\mu_{l-1})$-stable. As $T^{l-1}_{\mu}$ is closed for both $\beta_{l}$ and $\tau$ under $\mu_{l-1}$, all the $(\tau,\beta_{l})$-chains intersect $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$ are contained in $G[V(T^{l-1}_{\mu})]$ except for $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ which contains $y_{l}$ who is the only $(\tau,\beta_{l})$-exit of $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$, and all vertices in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$ are contained in $T^{l-1}_{\mu}$. Moreover, since $\beta_{l}\notin\{\tau,\beta_{l-1}\}$ by (13), $y_{l-1}\prec y_{l}$ by (12), and $T^{l}_{\mu}$ is elementary by $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$), we have $\beta_{l}\notin(\overline{\mu}_{l-1}(T^{l-1}_{\mu}(y_{l}))\cup\{\beta_{l-1}\}$. So $\beta_{l}\notin\mu_{l-1}(P\cup P_{ex}\cup D\cup P_{1})$ by $(C_{l-1})$ (when $l\geq 2$) and $(c_{i})$ (when $l=1$). Thus $\tau$ is not used on edges of $P\cup P_{ex}\cup D\cup P_{1}-G[V(T^{l-1}_{\mu})]$ and $\beta_{l}$ is not used on edges of $(P\cup P_{ex}\cup D\cup P_{1})\cap G[V(T^{l-1}_{\mu})]$ under $\mu^{\prime}$ and $\mu_{l}^{\prime}$. Therefore, since all vertices in $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$ are contained in $T^{l-1}_{\mu}$, none of the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ under $\mu_{l}^{\prime}$ is in the subpath of $P_{r}(\tau,\beta_{l},\mu_{l}^{\prime})$ from $r$ to $y_{l}$. Consequently, we have $(E_{l})$ as we switch all the $\beta_{l}$ colored edges of $P\cup P_{ex}\cup D\cup P_{1}$ not in $P_{r}(\pi,\beta_{l},\mu_{l}^{\prime})$ to $\tau$ by Kempe changes, and none of these $(\tau,\beta_{l})$-chains intersect $T^{l-1}_{\mu}$ under $\mu_{l}^{\prime}$. Furthermore,
@@
So we may assume that $l=1$. So $\beta_{0}=\eta=\delta_{h}$. In this case, $P_{y_{1}}(\beta_{1},\delta_{h},\mu^{\prime})$ is the same path as the $(\tau,\delta_{h})$-exit path of $T_{i}$, which is the subpath $P_{1}[y_{1},r^{\prime}]$ of $P_{1}[w,r^{\prime}]$. Since $T_{n+1}$ is $\operatorname{RE}$ finished under $\varphi$, all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T_{n+1})]$. Since $P_{1}[w,r^{\prime}]$ does not intersect $T_{n+1}$ by assumption (iv) of $P$, $P_{1}[y_{1},r^{\prime}]$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi$. Because all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\varphi$ are the same as under $\varphi_{i}$ by $(a_{k})$ for all $1\leq k\leq i$, we see that $P_{1}[y_{1},r^{\prime}]$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\varphi_{i}=\mu_{0}$. Thus all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\mu_{0}$ are the same as under $\mu_{1}^{\prime}$ by the construction of $\mu_{1}^{\prime}$. Since $\beta_{1}\neq\delta_{h}$ ($\beta_{0}=\delta_{h}$) and $\tau\notin S_{h}$ (see (13)), we see that all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ under $\mu_{1}$ are the same as under $\mu_{1}^{\prime}$. Thus we have $(A_{1})$ by (14).
By (A2), $(B_{l-1})$ (when $l\geq 2$), and $(b_{i})$ (when $l=1$) under $\mu_{l-1}$, we have
@@
By $(A_{l})$ and (15), we see that $T_{h}$ has the same $(\gamma_{h},\delta_{h})$-exit vertex under $\mu_{l}$ as under $\mu_{l-1}$ if $h\geq 1$. Recall that $y_{l}\notin T_{j-1}$ by (12). Since $\mu_{l}$ is $(T^{l-1}_{\mu}(y_{l})-y_{l},D_{j-1},\mu_{l-1})$-stable, $y_{l}\notin T_{j-1}$, and $T^{l-1}_{\mu}$ has the same rung under $\mu_{l-1}$ as $T_{j}$ under $\varphi$ by $(B_{j-1})$ (when $j\geq 2$) and $(b_{i})$ (when $j=1$), we see that $\mu_{l}$ is also $(T_{j-1},D_{j-1},\mu_{l-1})$-stable, and thus is a $T_{j-1}+f_{j-1}$ mod $\mu_{l-1}$ coloring by (A2.5). Therefore, following Corollary 3.1, $T^{l-1}_{\mu}(y_{l})$ is an ETT with the same (j1)(j-1)-rungs under $\mu_{l}$ as $T^{l-1}_{\mu}$ under $\mu_{l-1}$. Since $T^{i}$ is a closure of $T^{l-1}_{\mu}(y_{l})$ under $\mu_{i}$ and $j-1\leq n$, we have $(B_{l})$ by (A1) on $T^{l}_{\mu}$, $(B_{l-1})$ (when $l\geq 2$) and $(b_{i})$ (when $l=1$).
@@
Note that $T^{l}_{\mu}(x)=T_{j}(x)$ following $(A_{l})$ and $(B_{l})$. We claim that $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$. Otherwise, by $(C_{l})$, we have that $V(P_{1})\subseteq V(T^{l}_{\mu})$ by TAA. So $w\in V(T^{l}_{\mu})$. Because $T^{l}_{\mu}$ is elementary by $(B_{l})$, we have $\alpha_{i}\in\overline{\mu}_{l}(w)$ by $(D_{l})$ and the fact that $w\notin V(T_{j}(x))$. So $u$ and $v$, the two ends of $P$ are also in $T^{l}_{\mu}$ by $(C_{l})$ and TAA. However, this is a contradiction to $(D_{l})$ and the elementariness of $T^{l}_{\mu}$ and the fact that $u,v\notin V(T_{j}(x))$. So $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$, and therefore $T^{l}_{\mu}$ must have a $(\tau,\beta_{l})$-exit vertex $y_{l+1}$ in $P_{r}(\tau,\beta_{l},\mu_{l})$ by the elementariness of $T^{l}_{\mu}$. This proves the first half of $(F_{l})$.
We then claim that $P_{r}(\tau,\beta_{l},\mu_{l})$ contains a vertex in $T^{l}_{\mu}$ that is after $y_{l}$ along the direction from $r$. Indeed, if $w\notin T^{l}_{\mu}$, then the two edge disjoint paths $P_{1}[r,w]$ and $P_{1}[w,y_{1}]$ provide us two $\beta_{l}$ colored boundary edges of $T^{l}_{\mu}$ which are also contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$ by $(E_{l})$ and $(C_{l})$, and one of them provide a vertex in $T^{l}_{\mu}$ that is after $y_{l}$ along the direction from $r$. So we assume that $w\in T^{l}_{\mu}$. Then because $w\notin V(T_{j}(x))=V(T^{l}_{\mu})(x)$, we have $\alpha_{i}\in\overline{\mu}_{l}(w)$ by $(D_{l})$ and the elementariness of $T^{l}_{\mu}$ followed from $(B_{l})$. If one of $u$ and $v$, say $v$, is contained in $T^{l}_{\mu}$, then $\beta_{i}\in\overline{\mu}_{l}(v)$ by the elementariness of $T^{l}_{\mu}$, $(D_{l})$ and the fact that $u,v\notin V(T_{j}(x))$. So both $u$ and $v$ are contained in $T^{l}_{\mu}$ by TAA and $(C_{l})$, a contradiction to $(D_{l})$ and the elementariness of $T^{l}_{\mu}$. Thus none of $u$ and $v$ is contained in $T^{l}_{\mu}$, and in this case $P_{2}$ and $P_{3}$ (see (3)) provide us two $\beta_{l}$ colored boundary edges of $T^{l}_{\mu}$ which are also contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ after the vertex $y_{l}$ along the direction from $r$ by $(E_{l})$ and $(C_{l})$. Similarly as earlier, we have as claimed.
@@
With the six claims above and the construction algorithm of $\mu_{l}$ earlier, we see that we can always recursively construct $\mu_{l}$ with a new vertex $y_{l+1}$ with $x_{i}=y_{0}\prec y_{1}\prec...\prec y_{l+1}$ along $T^{l}_{\mu}$ by (12), unless $\alpha_{i}\in\overline{\mu}_{l}(T^{l}_{\mu}(y_{l+1}))$. But since we have a finite graph, at a certain point we have to stop at some $\mu_{l}$ with $\alpha_{i}\in\overline{\mu}_{l}(T^{l}_{\mu}(y_{l+1}))$. For convenience, we still denote the corresponding coloring by $\mu_{l}$ and use index $l$ for everything we established earlier. Note that $l=0$ may happen.
@@
By $(F_{l})$ (when $l\geq 1$) and (10) (when $l=0$), we see that $\beta_{l}\notin\overline{\mu}_{l}(T^{l}_{\mu})$. Since $u,v\notin V(T_{j}(x))$, following the elementariness of $T^{l}_{\mu}$, $(D_{l})$ (when $l\geq 1$) and $(d_{i})$ (when $l=0$), none of $u$ and $v$ is contained in $T^{l}_{\mu}$. Thus following TAA, (16), $(C_{l})$ (when $l\geq 1$) and $(c_{i})$ (when $l=0$), the paths $P_{2}$ and $P_{3}$ (see (3)) contain two $\beta_{l}$ colored boundary edges $f$ and $g$ of $T^{l}_{\mu}$, where $f\in E(P_{2})$ is incident with $u^{\prime}\in V(T^{l}_{\mu})$ and $g\in E(P_{3})$ is incident with $v^{\prime}\in V(T^{l}_{\mu})$, such that $P_{2}[u^{\prime},u]$ and $P_{3}[v^{\prime},v]$ do not share edges with $G[V(T^{l}_{\mu})]$. Note that all the $\beta_{l}$ colored edges of $P_{2}\cup P_{3}$ are contained in $P_{r}(\tau,\beta_{l},\mu_{l})$ by $(E_{l})$ (when $l\geq 1$), $(c_{i})$ and assumption (iv) of $P$ (when $l=0$). Therefore, since $y_{l+1}$ is the $(\tau,\beta_{l})$-exit vertex of $T^{l}_{\mu}$ in $P_{r}(\tau,\beta_{l},\mu_{l})$, we see that both $u^{\prime}$ and $v^{\prime}$ are in the subpath of $P_{r}(\tau,\beta_{l},\mu_{l})$ from $r$ to $y_{l+1}$. We then let $u^{*}$ (and $v^{*}$) be the nearest vertex in $P_{2}$ (and $P_{3}$) to $u$ (and $v$) shared by the subpath of $P_{r}(\tau,\beta_{l},\mu_{l})$ from $r$ to $y_{l+1}$ and $P_{2}[u^{\prime},u]$ (and $P_{3}[v^{\prime},v]$), respectively. Without loss of the generality, we may assume $u^{*}$ is before $v^{*}$ along $P_{r}(\tau,\beta_{l},\mu_{l})=P_{4}$ from $r$. Now $P_{5}=P_{4}[r,u^{*}]\cup P_{2}[u^{*},u]$ and $P_{6}=P_{4}[y_{l+1},v^{*}]\cup P_{3}[v^{*},v]$ are two edge disjoint paths from $r$ to $u$ and from $y_{l+1}$ to $v$, respectively. Moreover, by (13), (16), $(C_{l})$ (when $l\geq 1$) and $(c_{i})$ (when $l=0$), we have
(17) $\mu_{l}(P_{5}\cup P_{6})\subseteq(\overline{\mu}_{l}(T^{l}_{\mu}(y_{l}))\cup\{\beta_{l}\}-\{\alpha_{i}\})$.
@@
Finally, we consider the coloring $\phi=\mu_{l}/(G-T^{l}_{\mu},\beta_{l+1},\tau)$, where $\beta_{l+1}$ is an arbitrary color in $\overline{\mu}_{l}(y_{l+1})$. Note that $\beta_{l+1}=\alpha_{i}$ may happen. It is now routine to check that $T^{l}_{\mu}$ remains an ETT with the same (j1)(j-1)-rungs under $\phi$ as $T_{j}$ under $\varphi$. Moreover, $P_{y_{l+1}}(\beta_{l},\beta_{l+1},\phi)$ is the $(\beta_{l},\beta_{l+1})$-exit path of $T^{l}_{\mu}$. By (17), no matter whether $\beta_{l+1}=\alpha_{i}$ or not, by the elementariness of $T^{l}_{\mu}$ under $\mu_{l}$ and $\phi$, we see that $\beta_{l+1}\notin\mu_{l}(P_{5}\cup P_{6})$. Thus $\phi(P_{5}\cup P_{6})\subseteq(\overline{\phi}(T^{l}_{\mu}(y_{l+1}))\cup\{\beta_{l}\})$, $\beta_{l+1}\notin\phi(P_{5}\cup P_{6})\cap G[V(T^{l}_{\mu})])$, and $\tau\notin\phi(P_{5}\cup P_{6})-G[V(T^{l}_{\mu})])$. In the case that $\beta_{l+1}=\delta_{h}$, we see that all the $(\gamma_{h},\delta_{h})$-ears of $T_{h}$ are contained in $G[V(T^{l}_{\mu})]$ under $\phi$, so $P_{y_{l+1}}(\beta_{l},\beta_{l+1},\phi)$ does not intersect any $(\gamma_{h},\delta_{h})$-ear of $T_{h}$ under $\phi$. In addition, it is also routine to check that $\tau\notin(\{\beta_{l+1}\}\cup S_{h})$ and $(ii)$ in Lemma 4.5 holds for $P_{5}$ and $P_{6}$. So by applying Lemma 4.5 with $P_{y_{l+1}}(\beta_{l},\beta_{l+1},\phi)$ in place of $P_{ex}$, $r$ in place of $z$, $\beta_{l}$ in place of $\delta$, $\beta_{l+1}$ in place of $\alpha$, and $\tau$ in place of $\pi$, we have a contradiction between $(II)$ and $(III)$, because $P_{5}$ and $P_{6}$ have different ends $u$ and $v$.